# On the speed of coming down from infinity for (sub)critical branching processes with pairwise interactions

### v1

@@ 3.5 Proof of Theorem 4
$\mathbb{E}\Big{[}\Big{|}\int_{r}^{t}(\mathbf{m}_{\rm br}Z_{s}^{(\infty)}+\mathbf{m}_{\rm int}Z_{s}^{(\infty)}(Z_{s}^{(\infty)}-1)){\rm d}s\Big{|}\Big{]}<\infty.$ (75)
@@
$Z_{t}^{(\infty)}=Z_{r}^{(\infty)}+\int_{r}^{t}(\mathbf{m}_{\rm br}Z_{s}^{(\infty)}+\mathbf{m}_{\rm int}Z_{s}^{(\infty)}(Z_{s}^{(\infty)}-1)){\rm d}s+\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(\infty)},u,y,x)\tilde{N}({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$ (76)
@@
$\quad\quad\quad-\mathbf{m}_{\rm int}\int_{(r,t]}\int_{\mathcal{X}}sH(Z_{s-}^{(\infty)},u,y,x)\tilde{N}({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$ (77)
@@
Corollary 2 (ii) implies (75) and thus, Lemma 14 implies (76) (see e.g., [33, Theorem 8.23]). Integration by parts and (76) imply (15). ∎
Recall, from the beginning of Section 3.1, the definitions of the Poisson random measures $N_{\rm br}$ and $N_{\rm int}$ on +××$\mathbb{R}_{+}\times\mathbb{N}\times\mathbb{N}_{-1}$ and +×$\mathbb{R}_{+}\times\Delta\times\mathbb{N}_{-1}$ with intensity measures ${\rm d}s\otimes q_{\rm br}({\rm d}y,{\rm d}x)$ and ${\rm d}s\otimes q_{\rm br}({\rm d}y,{\rm d}x)$, respectively, where
@@
Lemma 16.
Suppose that $\mathbf{m}_{\rm int}<0$ and (${\bf H4}$) hold. Fix $0<t_{0}<T(\infty)$. The processes $(M_{t}^{\rm br})_{t\geq 0}$ and $(M_{t}^{\rm int})_{t\geq 0}$ given by
$M_{t}^{\rm br}:=-\mathbf{m}_{\rm int}\int_{0}^{t\wedge t_{0}}\int_{\mathbb{N}}\int_{\mathbb{N}_{-1}}sx\mathbf{1}_{\{0<y\leq Z_{s-}^{(\infty)}\}}\tilde{N}_{\rm br}({\rm d}s,{\rm d}y,{\rm d}x)$ (78)
@@
$M_{t}^{\rm int}:=-\mathbf{m}_{\rm int}\int_{0}^{t\wedge t_{0}}\int_{\Delta}\int_{\mathbb{N}_{-1}}sx\mathbf{1}_{\{y\in\Delta_{Z_{s-}^{(\infty)}}\}}\tilde{N}_{\rm int}({\rm d}s,{\rm d}y,{\rm d}x),\quad\text{for}\quad t\geq 0,$ (79)
@@
$\langle M^{\rm br}\rangle_{t}=(\mathbf{m}_{\rm int})^{2}\bm{\sigma}^{2}_{{\rm br}}\int_{0}^{t\wedge t^{\ast}}s^{2}Z_{s}^{(\infty)}{\rm d}s$ (80)
@@
$\langle M^{\rm int}\rangle_{t}=(\mathbf{m}_{\rm int})^{2}\bm{\sigma}^{2}_{{\rm int}}\int_{0}^{t\wedge t^{\ast}}s^{2}Z_{s}^{(\infty)}(Z_{s}^{(\infty)}-1){\rm d}s,\quad\text{for}\quad t\geq 0.$ (81)
@@
$\mathbb{E}\Big{[}\sup_{s\leq t}|M_{s}^{\rm int}|^{p}\Big{]}\leq C_{2}(p)t^{p/2},\quad\text{for}\quad t\leq T.$ (83)
@@
$\int_{0}^{t\wedge t^{\ast}}\int_{\mathbb{N}}\int_{\mathbb{N}_{-1}}s^{2}x^{2}\mathbf{1}_{\{0<y\leq Z_{s-}^{(\infty)}\}}q_{\rm br}({\rm d}y,{\rm d}x){\rm d}s$ $=\bm{\sigma}^{2}_{{\rm br}}\int_{0}^{t\wedge t^{\ast}}s^{2}Z_{s}^{(\infty)}{\rm d}s$
@@
$\int_{0}^{t\wedge t^{\ast}}\int_{\Delta}\int_{\mathbb{N}_{-1}}s^{2}x^{2}\mathbf{1}_{\{y\in\Delta_{Z_{s-}^{(\infty)}}\}}q_{\rm int}({\rm d}y,{\rm d}x){\rm d}s$ $=\bm{\sigma}^{2}_{{\rm int}}\int_{0}^{t\wedge t^{\ast}}s^{2}Z_{s}^{(\infty)}(Z_{s}^{(\infty)}-1){\rm d}s,\quad\text{for}\quad t\geq 0.$
By Corollary 2 (i) and (${\bf H4}$), for any $0\leq T<\infty$, there exist constants $0<C_{1},C_{2}<\infty$ such that
$\mathbb{E}\Big{[}(\mathbf{m}_{\rm int})^{2}\int_{0}^{t\wedge t^{\ast}}\int_{\mathbb{N}}\int_{\mathbb{N}_{-1}}s^{2}x^{2}\mathbf{1}_{\{0<y\leq Z_{s-}^{(\infty)}\}}q_{\rm br}({\rm d}y,{\rm d}x){\rm d}s\Big{]}\leq C_{1}t^{2}$ (84)
@@
$\mathbb{E}\Big{[}(\mathbf{m}_{\rm int})^{2}\int_{0}^{t\wedge t^{\ast}}\int_{\Delta}\int_{\mathbb{N}_{-1}}s^{2}x^{2}\mathbf{1}_{\{y\in\Delta_{Z_{s-}^{(\infty)}}\}}q_{\rm int}({\rm d}y,{\rm d}x){\rm d}s\Big{]}\leq C_{2}t,\quad\text{for}\quad t\leq T.$ (85)
It follows from standard properties of Poisson integration [33, Theorem 8.23] that $(M_{t}^{\rm br})_{t\geq 0}$ and $(M_{t}^{\rm int})_{t\geq 0}$ are well-defined locally square-integrable martingales with quadratic variation (80) and (81), respectively. For $p=2$, (82) and (83) are consequence of (84), (85) and Doob’s maximal inequality. For $p\in(0,2)$, (82) and (83) follow by Jensen’s inequality. This completes the proof. ∎
Lemma 17.
@@
$\mathbb{E}\Big{[}\sup_{s\leq t}\Big{|}-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1\Big{|}^{p}\Big{]}\leq C(p)(t^{p/2}\vee t^{p}),\quad\text{for}\quad t\leq T.$
@@
$-\int_{r}^{t}(\mathbf{m}_{\rm int}\mathbf{m}_{\rm br}sZ_{s}^{(\infty)}+\mathbf{m}_{\rm int}Z_{s}^{(\infty)}+(\mathbf{m}_{\rm int})^{2}sZ_{s}^{(\infty)}(Z_{s}^{(\infty)}-1)){\rm d}s$
@@
$-\int_{r}^{t}s^{-1}(-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)})(-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1){\rm d}s+\int_{0}^{t}(\mathbf{m}_{\rm int})^{2}sZ_{s}^{(\infty)}{\rm d}s-\int_{r}^{t}\mathbf{m}_{\rm int}\mathbf{m}_{\rm br}sZ_{s}^{(\infty)}{\rm d}s.$
Fix $0\leq T<T(\infty)$. Let $M_{t}=M_{t}^{\rm br}+M_{t}^{\rm int}$, for $t\geq 0$, where $(M_{t}^{\rm br})_{t\geq 0}$ and $(M_{t}^{\rm int})_{t\geq 0}$ are the martingales defined in (78) and (79), respectively, such that $T\leq t_{0}<T(\infty)$. By (15) and [6, Lemma 10] (with $g(s)=-\mathbf{m}_{\rm int}Z_{s}^{(\infty)}(-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1)$), we have that
$\sup_{r\leq s\leq t}\left|-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1\right|\leq 2\Big{(}\left|-\mathbf{m}_{\rm int}rZ_{r}^{(\infty)}-1\right|+|M_{r}|+\sup_{r\leq s\leq t}|M_{s}|+\int_{0}^{t}|(\mathbf{m}_{\rm int})^{2}-\mathbf{m}_{\rm int}\left|\mathbf{m}_{\rm br}|\right|sZ_{s}^{(\infty)}{\rm d}s\Big{)},$
for $r\leq t\leq T$. Then, Corollary 2 (i), (82) and (83) imply that there are constants $0<C_{1}(p),C_{2}(p)<\infty$ such that
$\mathbb{E}\Big{[}\sup_{r\leq s\leq t}\left|-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1\right|^{p}\Big{]}\leq 2\cdot 4^{p}\Big{(}\mathbb{E}\Big{[}\left|-\mathbf{m}_{\rm int}rZ_{r}^{(\infty)}-1\right|^{p}\Big{]}+\mathbb{E}[|M_{r}|^{p}]+C_{1}(p)t^{p/2}+C_{2}(p)t^{p}\Big{)}.$
Our claim follows by letting $r\downarrow 0$, and using Theorem 3 and the inequalities (82) and (83). ∎
Lemma 18.
Suppose that $\mathbf{m}_{\rm int}<0$ and (${\bf H4}$) hold. Fix $0<t_{0}<T(\infty)$. For all $t\geq 0$, the integral
$A_{t}=-\int_{0}^{t\wedge t_{0}}\Big{(}\mathbf{m}_{\rm int}\mathbf{m}_{\rm br}sZ_{s}^{(\infty)}+\mathbf{m}_{\rm int}Z_{s}^{(\infty)}+(\mathbf{m}_{\rm int})^{2}sZ_{s}^{(\infty)}(Z_{s}^{(\infty)}-1)\Big{)}{\rm d}s$
@@
$A_{t}=-\int_{0}^{t\wedge t_{0}}s^{-1}\Big{(}-s\mathbf{m}_{\rm int}Z_{s}^{(\infty)}-1\Big{)}{\rm d}s+U_{t},\quad\text{for}\quad t\geq 0,$
@@
$\mathbb{E}\Big{[}\sup_{s\leq t}|U_{s}|\Big{]}\leq C(t\vee t^{2}),\quad\text{for}\quad t\leq T.$ (86)
@@
$A_{t}$ $=-\int_{0}^{t\wedge t_{0}}s^{-1}(-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1)^{2}{\rm d}s-\int_{0}^{t\wedge t_{0}}s^{-1}(-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1){\rm d}s$
$\quad\quad\quad+\int_{0}^{t\wedge t_{0}}(\mathbf{m}_{\rm int})^{2}sZ_{s}^{(\infty)}{\rm d}s-\int_{0}^{t\wedge t_{0}}\mathbf{m}_{\rm int}\mathbf{m}_{\rm br}sZ_{s}^{(\infty)}{\rm d}s.$ (87)
Corollary 2 (i) and Lemma 17 readily imply that each of the integrals in (3.5) is a well-defined Lebesgue integral for all $t\geq 0$ simultaneously, almost surely. This proves the first claim.
@@
$U_{t}$ $=-\int_{0}^{t\wedge t_{0}}s^{-1}(-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1)^{2}{\rm d}s+\int_{0}^{t\wedge t_{0}}(\mathbf{m}_{\rm int})^{2}sZ_{s}^{(\infty)}{\rm d}s-\int_{0}^{t\wedge t_{0}}\mathbf{m}_{\rm int}\mathbf{m}_{\rm br}sZ_{s}^{(\infty)}{\rm d}s,\quad\text{for}\quad t\geq 0,$
we deduce (86) by using Corollary 2 (i) and Lemma 17. This completes the proof. ∎
Lemma 19.
Suppose that $\mathbf{m}_{\rm int}<0$ and (${\bf H4}$) hold. Let $(M_{t}^{\rm br})_{t\geq 0}$, $(M_{t}^{\rm int})_{t\geq 0}$ and $(U_{t})_{t\geq 0}$ be the processes defined in Lemma 16 and Lemma 18, respectively, for some $0<t_{0}<T(\infty)$. Then,
$-\mathbf{m}_{\rm int}tZ_{t}^{(\infty)}$ $=1-\int_{0}^{t}s^{-1}\Big{(}-s\mathbf{m}_{\rm int}Z_{s}^{(\infty)}-1\Big{)}{\rm d}s+M_{t}^{\rm br}+M_{t}^{\rm int}+U_{t},\quad 0\leq t\leq t_{0}.$ (88)
@@
By Lemma 16 and Lemma 18, the process introduced in (88) is well-defined. On the other hand, by (15), Lemma 16 and Lemma 18, we have that, for $0<r\leq t\leq t_{0}$,
$-\mathbf{m}_{\rm int}tZ_{t}^{(\infty)}=-\mathbf{m}_{\rm int}rZ_{r}^{(\infty)}+(A_{t}-A_{r})+(M_{t}^{\rm br}-M_{r}^{\rm br})+(M_{t}^{\rm int}-M_{r}^{\rm int}).$
First, Theorem 3 implies that $-\mathbf{m}_{\rm int}rZ_{r}^{(\infty)}$ converges to $1$ in $L_{2}$, as $r\downarrow 0$. Second, by (82) and (83), we have that $M_{r}^{\rm br}$ and $M_{r}^{\rm int}$ converge to 00 in $L_{2}$, as $r\downarrow 0$. Finally, (3.5), Corollary 2 (i) and Lemma 17 imply that there is a constant $0<C<\infty$ such that
$\mathbb{E}[|A_{r}|]\leq C(r^{1/2}\vee r\vee r^{2}),$
which implies that $A_{r}$ converges to 00 in $L_{1}$, as $r\downarrow 0$. The above implies that for any fixed $0<t\leq t_{0}$, as $r\downarrow 0$, both the left-hand side and the right-hand side of (15) converge in probability to the corresponding left and right hand side of (88). Due to the càdlàg property of all the processes under consideration, identity (88) holds for all $0\leq t\leq t^{\ast}$ simultaneously. ∎
Henceforth, we fix $0<t_{0}<\infty$. Let $(D_{t})_{t\geq 0}$ be the stochastic process given by
$D_{t}:=-\int_{0}^{t}s{\rm d}M_{s}^{\rm int},\quad\text{for}\quad t\geq 0.$
Clearly, $(D_{t})_{t\geq 0}$ is well-defined. In particular, by (79),
$D_{t}=\mathbf{m}_{\rm int}\int_{0}^{t\wedge t_{0}}\int_{\Delta}\int_{\mathbb{N}_{-1}}s^{2}x\mathbf{1}_{\{y\in\Delta_{Z_{s-}^{(\infty)}}\}}\tilde{N}_{\rm int}({\rm d}s,{\rm d}y,{\rm d}x),\quad\text{for}\quad t\geq 0.$
@@
$M_{k,t}^{\rm int}:=-\mathbf{m}_{\rm int}\int_{0}^{t\wedge t_{0}}\int_{\Delta}\int_{\mathbb{N}_{-1}}sx\mathbf{1}_{\{|x|\leq k\}}\mathbf{1}_{\{y\in\Delta_{Z_{s-}^{(\infty)}}\}}\tilde{N}_{\rm int}({\rm d}s,{\rm d}y,{\rm d}x),\quad\text{for}\quad t\geq 0.$
Note that $(M_{k,t}^{\rm int})_{t\geq 0}$ is a well-defined locally square-integrable martingale (see e.g., the argument used in the proof of Lemma 16). Then, define the stochastic process $(D_{k,t})_{t\geq 0}$ as follows
$D_{k,t}:=-\int_{0}^{t}s{\rm d}M_{k,t}^{\rm int}=\mathbf{m}_{\rm int}\int_{0}^{t\wedge t_{0}}\int_{\Delta}\int_{\mathbb{N}_{-1}}s^{2}x\mathbf{1}_{\{|x|\leq k\}}\mathbf{1}_{\{y\in\Delta_{Z_{s-}^{(\infty)}}\}}\tilde{N}_{\rm int}({\rm d}s,{\rm d}y,{\rm d}x),\quad\text{for}\quad t\geq 0.$ (89)
Lemma 20.

### v2

@@ 3.1 Integral representation of BPI processes
$\mathbb{E}\Big{[}\Big{|}\int_{r}^{t}(\mathbf{m}_{\rm br}Z_{s}^{(\infty)}+\mathbf{m}_{\rm int}Z_{s}^{(\infty)}(Z_{s}^{(\infty)}-1)){\rm d}s\Big{|}\Big{]}<\infty.$ (3.76)
@@
$Z_{t}^{(\infty)}$ $=Z_{r}^{(\infty)}+\int_{r}^{t}(\mathbf{m}_{\rm br}Z_{s}^{(\infty)}+\mathbf{m}_{\rm int}Z_{s}^{(\infty)}(Z_{s}^{(\infty)}-1)){\rm d}s$
$\quad\quad\quad+\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(\infty)},u,y,x)\tilde{N}({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$ (3.77)
@@
$\quad\quad\quad-\mathbf{m}_{\rm int}\int_{(r,t]}\int_{\mathcal{X}}sH(Z_{s-}^{(\infty)},u,y,x)\tilde{N}({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$ (3.78)
@@
Corollary 3.11 (ii) implies (3.76) and thus, Lemma 3.12 implies (3.13) (see e.g., Theorem 8.23 in [32]). Integration by parts and (3.13) imply (3.13). ∎
Recall, from the beginning of Section 3.1, the definitions of the Poisson random measures $N_{\rm br}$ and $N_{\rm int}$ on $\mathbb{R}_{+}\times\mathbb{N}\times\mathbb{N}_{-1}$ and $\mathbb{R}_{+}\times\Delta\times\mathbb{N}_{-1}$ with intensity measures ${\rm d}s\otimes q_{\rm br}({\rm d}y,{\rm d}x)$ and ${\rm d}s\otimes q_{\rm br}({\rm d}y,{\rm d}x)$, respectively, where
@@
Lemma 3.14.
Suppose that $\mathbf{m}_{\rm int}<0$ and ($\bf H4$}) hold. Fix $0<t^{\ast}<T(\infty)$. The processes $(M_{t}^{\rm br})_{t\geq 0}$ and $(M_{t}^{\rm int})_{t\geq 0}$ given by
$M_{t}^{\rm br}:=-\mathbf{m}_{\rm int}\int_{0}^{t\wedge t^{\ast}}\int_{\mathbb{N}}\int_{\mathbb{N}_{-1}}sx\mathbf{1}_{\{0<y\leq Z_{s-}^{(\infty)}\}}\tilde{N}_{\rm br}({\rm d}s,{\rm d}y,{\rm d}x)$ (3.79)
@@
$M_{t}^{\rm int}:=-\mathbf{m}_{\rm int}\int_{0}^{t\wedge t^{\ast}}\int_{\Delta}\int_{\mathbb{N}_{-1}}sx\mathbf{1}_{\{y\in\Delta_{Z_{s-}^{(\infty)}}\}}\tilde{N}_{\rm int}({\rm d}s,{\rm d}y,{\rm d}x),\quad\text{for}\quad t\geq 0,$ (3.80)
@@
$\langle M^{\rm br}\rangle_{t}=(\mathbf{m}_{\rm int})^{2}\bm{\sigma}^{2}_{{\rm br}}\int_{0}^{t\wedge t^{\ast}}s^{2}Z_{s}^{(\infty)}{\rm d}s$ (3.81)
@@
$\langle M^{\rm int}\rangle_{t}=(\mathbf{m}_{\rm int})^{2}\bm{\sigma}^{2}_{{\rm int}}\int_{0}^{t\wedge t^{\ast}}s^{2}Z_{s}^{(\infty)}(Z_{s}^{(\infty)}-1){\rm d}s,\quad\text{for}\quad t\geq 0.$ (3.82)
@@
$\mathbb{E}\Big{[}\sup_{s\leq t}|M_{s}^{\rm int}|^{p}\Big{]}\leq C_{2}(p)t^{p/2},\quad\text{for}\quad t\leq T.$ (3.84)
@@
$\int_{0}^{t\wedge t^{\ast}}\int_{\mathbb{N}}\int_{\mathbb{N}_{-1}}s^{2}x^{2}\mathbf{1}_{\{0<y\leq Z_{s-}^{(\infty)}\}}q_{\rm br}({\rm d}y,{\rm d}x){\rm d}s$ $=\bm{\sigma}^{2}_{{\rm br}}\int_{0}^{t\wedge t^{\ast}}s^{2}Z_{s}^{(\infty)}{\rm d}s$ (3.85)
@@
$\int_{0}^{t\wedge t^{\ast}}\int_{\Delta}\int_{\mathbb{N}_{-1}}s^{2}x^{2}\mathbf{1}_{\{y\in\Delta_{Z_{s-}^{(\infty)}}\}}q_{\rm int}({\rm d}y,{\rm d}x){\rm d}s$ $=\bm{\sigma}^{2}_{{\rm int}}\int_{0}^{t\wedge t^{\ast}}s^{2}Z_{s}^{(\infty)}(Z_{s}^{(\infty)}-1){\rm d}s,\quad\text{for}\quad t\geq 0.$ (3.86)
By Corollary 3.11 (i) and ($\bf H4$}), for any $0\leq T<\infty$, there exist constants $0<C_{1},C_{2}<\infty$ such that
$\mathbb{E}\Big{[}(\mathbf{m}_{\rm int})^{2}\int_{0}^{t\wedge t^{\ast}}\int_{\mathbb{N}}\int_{\mathbb{N}_{-1}}s^{2}x^{2}\mathbf{1}_{\{0<y\leq Z_{s-}^{(\infty)}\}}q_{\rm br}({\rm d}y,{\rm d}x){\rm d}s\Big{]}\leq C_{1}t^{2}$ (3.87)
@@
$\mathbb{E}\Big{[}(\mathbf{m}_{\rm int})^{2}\int_{0}^{t\wedge t^{\ast}}\int_{\Delta}\int_{\mathbb{N}_{-1}}s^{2}x^{2}\mathbf{1}_{\{y\in\Delta_{Z_{s-}^{(\infty)}}\}}q_{\rm int}({\rm d}y,{\rm d}x){\rm d}s\Big{]}\leq C_{2}t,\quad\text{for}\quad t\leq T.$ (3.88)
It follows from standard properties of Poisson integration [32, Theorem 8.23] that $(M_{t}^{\rm br})_{t\geq 0}$ and $(M_{t}^{\rm int})_{t\geq 0}$ are well-defined locally square-integrable martingales with quadratic variation (3.81) and (3.82), respectively. For $p=2$, (3.83) and (3.84) are consequence of (3.87), (3.88) and Doob’s maximal inequality. For $p\in(0,2)$, (3.83) and (3.84) follow by Jensen’s inequality. This completes the proof. ∎
Lemma 3.15.
@@
$\mathbb{E}\Big{[}\sup_{s\leq t}\Big{|}-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1\Big{|}^{p}\Big{]}\leq C(p)(t^{p/2}\vee t^{p}),\quad\text{for}\quad t\leq T.$ (3.89)
@@
$-\int_{r}^{t}(\mathbf{m}_{\rm int}\mathbf{m}_{\rm br}sZ_{s}^{(\infty)}+\mathbf{m}_{\rm int}Z_{s}^{(\infty)}+(\mathbf{m}_{\rm int})^{2}sZ_{s}^{(\infty)}(Z_{s}^{(\infty)}-1)){\rm d}s$ (3.90)
@@
$-\int_{r}^{t}s^{-1}(-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)})(-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1){\rm d}s+\int_{r}^{t}(\mathbf{m}_{\rm int})^{2}sZ_{s}^{(\infty)}{\rm d}s-\int_{r}^{t}\mathbf{m}_{\rm int}\mathbf{m}_{\rm br}sZ_{s}^{(\infty)}{\rm d}s.$ (3.91)
Fix $0\leq T<T(\infty)$. Let $M_{t}=M_{t}^{\rm br}+M_{t}^{\rm int}$, for $t\geq 0$, where $(M_{t}^{\rm br})_{t\geq 0}$ and $(M_{t}^{\rm int})_{t\geq 0}$ are the martingales defined in (3.79) and (3.80), respectively, such that $T\leq t^{\ast}<T(\infty)$. By (3.13) and Lemma 10 in [6] (with $g(s)=-\mathbf{m}_{\rm int}Z_{s}^{(\infty)}(-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1)$), we have that
$\sup_{r\leq s\leq t}\left|-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1\right|$
$\quad\quad\quad\leq 2\Big{(}\left|-\mathbf{m}_{\rm int}rZ_{r}^{(\infty)}-1\right|+|M_{r}|+\sup_{r\leq s\leq t}|M_{s}|+\int_{0}^{t}|(\mathbf{m}_{\rm int})^{2}-\mathbf{m}_{\rm int}\mathbf{m}_{\rm br}|sZ_{s}^{(\infty)}{\rm d}s\Big{)},$ (3.92)
for $r\leq t\leq T$. Then, Corollary 3.11 (i), (3.83) and (3.84) imply that there are constants $0<C_{1}(p),C_{2}(p)<\infty$ such that
$\mathbb{E}\Big{[}\sup_{r\leq s\leq t}\left|-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1\right|^{p}\Big{]}$
$\quad\quad\quad\leq 2\cdot 4^{p}\Big{(}\mathbb{E}\Big{[}\left|-\mathbf{m}_{\rm int}rZ_{r}^{(\infty)}-1\right|^{p}\Big{]}+\mathbb{E}[|M_{r}|^{p}]+C_{1}(p)t^{p/2}+C_{2}(p)t^{p}\Big{)}.$ (3.93)
Our claim follows by letting $r\downarrow 0$, and using Theorem 2.2 and the inequalities (3.83) and (3.84). ∎
Lemma 3.16.
Suppose that $\mathbf{m}_{\rm int}<0$ and ($\bf H4$}) hold. Fix $0<t^{\ast}<T(\infty)$. For all $t\geq 0$, the integral
$A_{t}=-\int_{0}^{t\wedge t^{\ast}}\Big{(}\mathbf{m}_{\rm int}\mathbf{m}_{\rm br}sZ_{s}^{(\infty)}+\mathbf{m}_{\rm int}Z_{s}^{(\infty)}+(\mathbf{m}_{\rm int})^{2}sZ_{s}^{(\infty)}(Z_{s}^{(\infty)}-1)\Big{)}{\rm d}s$ (3.94)
@@
$A_{t}=-\int_{0}^{t\wedge t^{\ast}}s^{-1}\Big{(}-s\mathbf{m}_{\rm int}Z_{s}^{(\infty)}-1\Big{)}{\rm d}s+U_{t},\quad\text{for}\quad t\geq 0,$ (3.95)
@@
$\mathbb{E}\Big{[}\sup_{s\leq t}|U_{s}|\Big{]}\leq C(t\vee t^{2}),\quad\text{for}\quad t\leq T.$ (3.96)
@@
$A_{t}$ $=-\int_{0}^{t\wedge t^{\ast}}s^{-1}(-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1)^{2}{\rm d}s-\int_{0}^{t\wedge t^{\ast}}s^{-1}(-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1){\rm d}s$
$\quad\quad\quad+\int_{0}^{t\wedge t^{\ast}}(\mathbf{m}_{\rm int})^{2}sZ_{s}^{(\infty)}{\rm d}s-\int_{0}^{t\wedge t^{\ast}}\mathbf{m}_{\rm int}\mathbf{m}_{\rm br}sZ_{s}^{(\infty)}{\rm d}s.$ (3.97)
Corollary 3.11 (i) and Lemma 3.15 readily imply that each of the integrals in (3.4) is a well-defined Lebesgue integral for all $t\geq 0$ simultaneously, almost surely. This proves the first claim.
@@
$U_{t}$ $=-\int_{0}^{t\wedge t^{\ast}}s^{-1}(-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1)^{2}{\rm d}s+\int_{0}^{t\wedge t^{\ast}}(\mathbf{m}_{\rm int})^{2}sZ_{s}^{(\infty)}{\rm d}s-\int_{0}^{t\wedge t^{\ast}}\mathbf{m}_{\rm int}\mathbf{m}_{\rm br}sZ_{s}^{(\infty)}{\rm d}s,$ (3.98)
for $t\geq 0$, we deduce (3.96) by using Corollary 3.11 (i) and Lemma 3.15. This completes the proof. ∎
Lemma 3.17.
Suppose that $\mathbf{m}_{\rm int}<0$ and ($\bf H4$}) hold. Let $(M_{t}^{\rm br})_{t\geq 0}$, $(M_{t}^{\rm int})_{t\geq 0}$ and $(U_{t})_{t\geq 0}$ be the processes defined in Lemma 3.14 and Lemma 3.16, respectively, for some $0<t^{\ast}<T(\infty)$. Then,
$-\mathbf{m}_{\rm int}tZ_{t}^{(\infty)}$ $=1-\int_{0}^{t}s^{-1}\Big{(}-s\mathbf{m}_{\rm int}Z_{s}^{(\infty)}-1\Big{)}{\rm d}s+M_{t}^{\rm br}+M_{t}^{\rm int}+U_{t},\quad 0\leq t\leq t^{\ast}.$ (3.99)
@@
By Lemma 3.14 and Lemma 3.16, the process introduced in (3.99) is well-defined. On the other hand, by (3.13), Lemma 3.14 and Lemma 3.16, we have that, for $0<r\leq t\leq t^{\ast}$,
$-\mathbf{m}_{\rm int}tZ_{t}^{(\infty)}=-\mathbf{m}_{\rm int}rZ_{r}^{(\infty)}+(A_{t}-A_{r})+(M_{t}^{\rm br}-M_{r}^{\rm br})+(M_{t}^{\rm int}-M_{r}^{\rm int}).$ (3.100)
First, Theorem 2.2 implies that $\mathbf{m}_{\rm int}rZ_{r}^{(\infty)}$ converges to $1$ in $L_{2}$, as $r\downarrow 0$. Second, by (3.83) and (3.84), we have that $M_{r}^{\rm br}$ and $M_{r}^{\rm int}$ converge to 0 in $L_{2}$, as $r\downarrow 0$. Finally, (3.4), Corollary 3.11 (i) and Lemma 3.15 imply that there is a constant $0<C<\infty$ such that
$\mathbb{E}[|A_{r}|]\leq C(r^{1/2}\vee r\vee r^{2}),$ (3.101)
which implies that $A_{r}$ converges to 0 in $L_{1}$, as $r\downarrow 0$. The above implies that for any fixed $0<t\leq t^{\ast}$, as $r\downarrow 0$, both the left-hand side and the right-hand side of (3.13) converge in probability to the corresponding left and right hand side of (3.99). Due to the càdlàg property of all the processes under consideration, identity (3.99) holds for all $0\leq t\leq t^{\ast}$ simultaneously. ∎
Henceforth, we fix $0<t^{\ast}<\infty$. Let $(D_{t})_{t\geq 0}$ be the stochastic process given by
$D_{t}:=-\int_{0}^{t}s{\rm d}M_{s}^{\rm int},\quad\text{for}\quad t\geq 0.$ (3.102)
Clearly, $(D_{t})_{t\geq 0}$ is well-defined. In particular, by (3.80),
$D_{t}=\mathbf{m}_{\rm int}\int_{0}^{t\wedge t^{\ast}}\int_{\Delta}\int_{\mathbb{N}_{-1}}s^{2}x\mathbf{1}_{\{y\in\Delta_{Z_{s-}^{(\infty)}}\}}\tilde{N}_{\rm int}({\rm d}s,{\rm d}y,{\rm d}x),\quad\text{for}\quad t\geq 0.$ (3.103)
@@
$M_{k,t}^{\rm int}:=-\mathbf{m}_{\rm int}\int_{0}^{t\wedge t^{\ast}}\int_{\Delta}\int_{\mathbb{N}_{-1}}sx\mathbf{1}_{\{|x|\leq k\}}\mathbf{1}_{\{y\in\Delta_{Z_{s-}^{(\infty)}}\}}\tilde{N}_{\rm int}({\rm d}s,{\rm d}y,{\rm d}x),\quad\text{for}\quad t\geq 0.$ (3.104)
Note that $(M_{k,t}^{\rm int})_{t\geq 0}$ is a well-defined locally square-integrable martingale (see e.g., the argument used in the proof of Lemma 3.14). Then, define the stochastic process $(D_{k,t})_{t\geq 0}$ as follows
$D_{k,t}:=-\int_{0}^{t}s{\rm d}M_{k,t}^{\rm int}=\mathbf{m}_{\rm int}\int_{0}^{t\wedge t^{\ast}}\int_{\Delta}\int_{\mathbb{N}_{-1}}s^{2}x\mathbf{1}_{\{|x|\leq k\}}\mathbf{1}_{\{y\in\Delta_{Z_{s-}^{(\infty)}}\}}\tilde{N}_{\rm int}({\rm d}s,{\rm d}y,{\rm d}x),\quad\text{for}\quad t\geq 0.$ (3.105)
Lemma 3.18.